OpenStudy (anonymous):

If ƒ = {(2, 5), (3, 2) (4, 6), (5, 1), (7, 2)}, then ƒ is a function. True False

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the first numbers are all different so TRUE

OpenStudy (anonymous):

If ƒ (x ) = 4x - 3, what is ƒ (x )^-1?

OpenStudy (misty1212):

\[4y-3=x\\ 4y=x+2\\ y=\frac{x+3}{4}\\ f^{-1}=\frac{x+3}{4}\]

OpenStudy (anonymous):

If ƒ (x ) = 2x^2 + 3, find ƒ (3). 31 15 21 39

OpenStudy (misty1212):

\[f(3)=2\times 3^2+3\] and a calculator or \(18+3=21\)

OpenStudy (anonymous):

If ƒ = {(2, 3), (5, 7), (3, 3), (5, 4), (9, 1)}, what is the range? {2, 5, 3, 9} {3, 7, 3, 4, 1} all whole numbers {1, 3, 4, 7}

OpenStudy (misty1212):

the list of the second numbers \[\{1,3,4,7\}\]